OpenStudy (anonymous):

The graph of the function f(x) = ax^2 +xa + c passes through the point (-2,9) and (2,13). Find the values of a and c. Show working out thanks.

OpenStudy (anonymous):

it passes through the points (-2,9) and (2,13)\[f(-2) = 9\]\[f(2) = 13\]and then solve for a and c

hero (hero):

If you substitute the given values of (x, f(x)) into the given equation and create two systems, you'd be able to find a and c

hero (hero):

Well create two equations in one system.

OpenStudy (anonymous):

\[f(-2) = 9\]\[a(-2)^2 + a(-2) + c = 9\]\[2a + c = 9\]\[f(2) = 13\]\[a(2)^2 + a(2) + c = 13\]\[6a + c = 13\]

OpenStudy (experimentx):

Oo.. it thought the second coff was b

hero (hero):

It happens. I used to get caught by stuff like that myself.

hero (hero):

I almost made that same mistake when I first looked at it.

OpenStudy (experimentx):

I make lots of observational error ... I don't know why.

hero (hero):

I as well. You just have to plow through it and train your eyes.

OpenStudy (experimentx):

i guess so..